# Randomized block Krylov method for approximation of truncated tensor SVD

### v2

@@ 6 Conclusion and future works
In this paper, we introduced the randomized block Krylov subspace method for approximating the truncated tensor singular value decomposition (T-SVD). We have examined this problem from a theoretical perspective and conducted numerical experiments using both synthetic and real-world datasets to validate our theoretical findings. Our numerical results clearly showed that the proposed algorithm is not only fast but also capable of delivering accurate approximations of the truncated T-SVD efficiently. We also present two applications of this randomized algorithm in the fields of data compression and completion. Currently, we are exploring the use of the randomized block Krylov subspace method for other types of tensor decompositions.
@@ 7 Acknowledgment
The authors would like to thank the editor and three reviewers for constructive and useful comment, which improved the quality of the paper significantly.

### v3

@@ 6 Conclusion and future works
In this paper, we introduced the randomized block Krylov subspace method for approximating the truncated tensor singular value decomposition. We have currently examined this problem from a theoretical perspective and conducted numerical experiments using both synthetic and real-world datasets to validate our theoretical findings. Our numerical results clearly showed that the proposed algorithm is not only fast but also capable of delivering accurate approximations of the truncated T-SVD efficiently. We also present two applications of this randomized algorithm in the fields of data compression and completion. Currently, we are exploring the use of the randomized block Krylov subspace method for other types of tensor decompositions. A promising direction for future work is the application of our randomized block Krylov method to enhance the adversarial robustness of deep neural networks, a topic we are currently exploring.
@@ 7 Acknowledgment
The work was supported by the Ministry of Economic Development of the Russian Federation under Agreement No. 139-10-2025-034 dd. 19.06.2025, IGK  [^00] C313925P4D0002.